OpenStudy (anonymous):

Deanne is four more than three times older than Ted. The sum of their ages is 60. How old is Ted?

OpenStudy (nikato):

Let D=Deanne and T=Ted So D=4+3T D+T=60 Use substitution

OpenStudy (nikato):

Subtract T on both sides in the second equation and get D=60-T Substitute this into the first equation and get 60-T=4+3T Now solve for T